# SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors

### v3

@@ 3.3 Evaluation Metrics
In ML and CL, the outputs are simple numerical values or formatted strings and contain no error or warning message, we employ exact string matching to measure correctness.
@@
For RL, we employ different evaluation methods for different tasks. For structured C repositories, we use exact character matching to compare outputs. For Sudoku and 24-point problems, we use edit distance to compare results. For other types of repositories, we apply the Ratcliff/Obershelp (Ratcliff and Metzener, 1988) algorithm, which measures the similarity between two sequences by finding the longest common subsequence and computing a similarity ratio.
@@
For SC and TC, various tasks necessitate distinct evaluation methods. Specifically, (1) numerical simulations are evaluated using the average Relative Absolute Error (RAE), a widely used metric that measures the deviation between the estimated values and the ground truth values; (2) position-based tasks, such as binary search, are assessed through exact string matching; and (3) sorting tasks are evaluated by the rank correlation coefficient (Spearman, 1904), which quantifies the similarity in the ordering of elements.
For BG, where outputs contain error messages, we use the Jaccard similarity (Jaccard, 1901) between predicted and ground truth error messages. Jaccard similarity measures the overlap between two sets and is defined as , where and represent the sets of words in the predicted and ground truth error messages respectively. Jaccard similarity well captures keywords in strings, thus suitable for error message comparison.
@@
For DR, since the same code can produce different outputs in varying settings, which sometimes include warnings or errors, we again utilize Jaccard similarity. This metric appropriately handles cases where the ground truth is not an error message.
@@
For FL, the results consist of two parts: (1) whether the proof passes or not, and (2) if it fails, the associated error message. The evaluation proceeds as follows. If both predicted and ground truth results indicate a successful proof, the prediction is considered correct. If one is passed and the other is not, the prediction is incorrect. If both fail, we evaluate the accuracy of the error message. The error message consists of a list containing the error locations and descriptions. We compute the score of a prediction as , where is the number of errors in the ground truth, is the set of predicted error positions, represents the -th ground truth error position, represents the predicted error position corresponding to , is the indicator function which equals to when there exists and equals to when there’s not, is the ground truth error message for position , is the predicted error message for position , and is the Jaccard similarity function.
@@ 3.4 Dataset Statistics
Except for RL, which has 60 questions, and DR, which has 200 questions, all other subsets have 150 questions. Our dataset contains a total of 1,160 questions. Table 1 presents detailed statistics of SURGE, including the construction methods, problem sources, dataset quantities, number of examples in few-shot scenarios, evaluation metrics, and classification criteria for each subset.
@@
4 Experiments
4.1 Setup
Models.
We tested SURGE on open-source and closed-source models of different sizes, including both chat models and code models. The closed-source models include GPT-4o (2024-08-06) (OpenAI, 2024b), GPT-4o-mini (2024-07-18) (OpenAI, 2024a), Claude-3.5-Sonnet (2024-10-22) (Anthropic, 2024a), and Qwen-Max (2025-01-25) (Team, 2024b). The open-source models include LLaMA-3.1-{8, 70}B-Instruct555https://huggingface.co/collections/meta-llama/llama-31-669fc079a0c406a149a5738f, LLaMA-3.3-70B-Instruct666https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Qwen-2.5-{0.5, 1.5, 3, 7, 14, 32, 72}B-Instruct777https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, Qwen-2.5-Coder-{0.5, 1.5, 3, 7, 14, 32}B-Instruct888https://huggingface.co/collections/Qwen/qwen25-coder-66eaa22e6f99801bf65b0c2f and DeepSeek-V3999https://huggingface.co/deepseek-ai/DeepSeek-V3 (671B).
Settings.
We tested the above models on SURGE under 3 settings: 0-shot w/o CoT, 0-shot w/ CoT, and few-shot w/ CoT. CoT here means whether we use Chain-of-Thought (Wei et al., 2022) prompting, allowing the models to think step by step, or ask the models to answer directly. We set the temperature to , i.e. employing greedy decoding.
4.2 Results
Table 2 presents the performance of 21 models across 8 sub-datasets of SURGE under 3 different settings. The best performance for each sub-dataset in each setting is highlighted in bold. In this table, we provide a detailed breakdown of the models’ performance on the ML and BG sub-datasets.
From the experimental results, several notable findings emerge:
SURGE demonstrates strong discriminative ability, especially in certain subsets. Even the strongest models perform only moderately well, highlighting the value of our benchmark. The models exhibit significant performance differences across different subsets, reflecting the comprehensiveness of our dataset. Additionally, models that perform well in other tasks (e.g. code generation), such as Claude-3.5-Sonnet, also achieve substantial overall results on SURGE. This demonstrates the reasonableness of our dataset and its effectiveness in benchmarking LLMs as general-purpose surrogate code executors.
Chat models and coder models exhibit different performance patterns and are affected differently by prompting strategies. We observed that for chat and code models of the same size (e.g. Qwen-2.5-32B-Instruct and Qwen-2.5-Coder-32B-Instruct), code models outperform chat models in the zero-shot setting SURGE. However, in the other two settings, chat models perform better. This suggests that code models have stronger zero-shot surrogate capabilities, whereas chat models excel in reasoning and imitation abilities.
5 Analysis
| Model | Compiler | Standard | Optimization | ||||||
| GPT-4o | 14.75 | 14.21 | 22.08 | 15.33 | 14.61 | 20.02 | 7.66 | 1.63 | 1.22 |
5.1 The Impact of Language Type
In the ML subset, models perform best in Python, followed by C++. Python’s simple syntax and dynamic typing make it easier to process, while C++ benefits from its strong presence in competitive programming and system-level tasks. Models also perform well in Julia, likely due to its clean syntax and similarity to Python. However, performance drops significantly in Rust, where strict ownership and lifetime rules introduce complexity, making code harder to predict.
In the BG subset, when predicting the output of buggy code, models excel in Python but struggle with C++. Python’s clear error messages and interpreted nature aid prediction, whereas C++’s static typing, manual memory management, and potential for undefined behavior make error handling more difficult.
5.2 The Impact of Problem Difficulty
| Difficulty | 1 | 2 | 3 | 4 | 5 |
| Zero-shot | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Zero-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Few-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
5.3 The Impact of Program’s CPU Execution Time
We explore the relationship between the execution time of the given code through running with CPUs and the accuracy of using large language models as surrogate models to acquire the output. As introduced in Section 3.1, we recorded the CPU execution times of the programs in TC. We categorize problems in TC into distinct bins based on their execution times. Next, we calculated the average accuracy for each model across all samples within the same bin, as shown in Figure 4. We observed the trend of prediction accuracy of the model falling as the actual execution time required for the corresponding program prolonged. It’s especially worth noting that for computational tasks that require execution time longer than 1 second, even state-of-the-art models struggle to obtain even one correct answer.
5.4 The Impact of Variable Type
The DR subset in SURGE examines model performance in predicting code behavior under different environmental factors. Specifically, DR includes variations in C++ compiler version, C++ standard, and compilation optimization settings. Table 3 detailedly presents model performance across different prompting strategies in DR.
5.5 Error Analysis
To further understand the model’s performance and limitations regarding serving as general surrogate models, we categorize the Errors made by Claude-3.5-Sonnet, GPT-4o and Llama-3.1-8B-Instruct and Llama-3.1-70B-Instructin the CL subset of SURGE.
We first combined machine-assisted annotation and manual verification to identify most typical error types: 1. Code Language Knowledge: evaluating foundational programming language proficiency, 2. Context Awareness: measuring understanding of long text and repository-level code, 3. Conceptual Understanding: assessing comprehension of programming concepts, 4. Calculation Process: verifying computational step accuracy, 5. Code Understanding: testing comprehension of code logic and structure, 6. Data Interpretation: evaluating data processing and analysis capabilities, and 7. Calculation Accuracy: measuring precision in scientific computations. We then use LLM to label errors in the model’s responses. The categorized error statistics are demonstrated in Figure 5.
As models prompted with CoT exhibit fewer instances of most error types compared to zero-shot, especially for the Code Understanding capability of Llama, which shows a significant improvement over zero-shot, CoT prompting results in a universal performance gain on the overall performance. The main error types are Code Understanding, Calculation Process, and Calculation Accuracy. For zero-shot, the primary error is accuracy, but for CoT, the most frequent error is Calculation Process. This suggests that CoT can better grasp the overall code logic and produce more correct results, but it may still make mistakes in the chain of thought process. In general, CoT has fewer and smaller errors. From the model perspective, Llama has a clear lead in Conceptual Understanding errors, indicating its weaker ability to understand concepts. The Claude model has the fewest errors, showing better performance under our criteria. Moreover, the error distribution of Claude and GPT is quite similar, which may suggest they share a similar way of thinking.
5.6 Training Scale Analysis
Apart from benchmarking, we systematically investigate whether training scaling can affect LLMs’ surrogate execution capabilities, we conducted comprehensive scaling experiments by training models of varying sizes on different amounts of training data. We conducted comprehensive scaling experiments with a dataset of 260K formal proofs and a fixed training batch size of 128.
As demonstrated in Figure 2, both model size and training steps are crucial factors in determining surrogate execution accuracy. As we scale from 0.5B to 7B parameters, models consistently show improved learning efficiency and higher performance ceilings throughout the training process. Larger models not only learn faster in the early stages but also continue to improve for longer before plateauing, suggesting better utilization of the training data.
These empirical observations align with established scaling laws in language modeling, indicating that surrogate execution capabilities follow similar scaling patterns as other language tasks. The consistent performance improvements with increased model size and training steps suggest that larger language models have significant potential to evolve into general-purpose surrogate code executors.
6 Conclusion
In this paper, we introduce SURGE, a holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. The curated dataset spans multiple domains, bridging theoretical tasks with real-world applications. Through extensive empirical study, we argue that while current large language models possess the power to predict code execution results to a certain extent, there remains significant room for further improvements on grounding LLMs to facilitate general surrogate model.
Acknowledgements
This work was independently conducted by the authors and is self-funded by BHL without institutional affiliation.

### v4

@@ 3.3 Evaluation Metrics
In ML and CL, the outputs are simple numerical values or formatted strings, we employ exact string matching to measure correctness.
@@
For RL, we employ different evaluation methods for different tasks. For structured C repositories, we use exact character matching to compare outputs. For Sudoku and 24-point problems, we use edit distance to compare results. For other types of repositories, we apply the Ratcliff/Obershelp (Ratcliff and Metzener, 1988) algorithm.
@@
For SC and TC, various tasks necessitate distinct evaluation methods. Specifically, (1) numerical simulations are evaluated using the average Relative Absolute Error (RAE); (2) position-based tasks, such as binary search, are assessed through exact string matching; and (3) sorting tasks are evaluated by the rank correlation coefficient (Spearman, 1904). Details regarding the evaluation metrics can be found in Appendix C.4.2 and Appendix C.5.2.
@@
For BG, we use the Jaccard similarity (Jaccard, 1901) between predicted and ground truth error messages. For DR, since the same code can produce different outputs in varying settings, which sometimes include warnings or errors, we again utilize Jaccard similarity.
@@
For FL, the results consist of two parts: (1) whether the proof passes or not, and (2) if it fails, we evaluate the accuracy of the error message. The error message consists of a list containing the error locations and descriptions. We compute the score of a prediction as , where is the number of errors in the ground truth, is the set of predicted error positions, represents the -th ground truth error position, represents the predicted error position corresponding to , is the indicator function which equals to only when there exists , is the ground truth error message for position , is the predicted error message for position , and is the Jaccard similarity function.
| Model | ML | CL | RL | SC | TC | BG | DR | FL | Avg. | |||||||
| CPP | Rust | Python | Julia | Java | Others | CPP | Java | Python | ||||||||
| Zero-shot | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| Zero-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
| Few-shot Chain-of-Thought | ||||||||||||||||
| Claude-3.5-Sonnet | ||||||||||||||||
| DeepSeek-V3 | ||||||||||||||||
| GPT-4o | ||||||||||||||||
| Qwen-Max | ||||||||||||||||
| Qwen-2.5-7B-Instruct | ||||||||||||||||
| Qwen-2.5-32B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-7B-Instruct | ||||||||||||||||
| Qwen-2.5-Coder-32B-Instruct | ||||||||||||||||
| LLaMA-3.1-8B-Instruct | ||||||||||||||||
| LLaMA-3.1-70B-Instruct | ||||||||||||||||
@@ 3.4 Dataset Statistics
Table 1 presents detailed statistics of SURGE, including the construction methods, problem sources, dataset quantities, number of examples in few-shot scenarios, evaluation metrics, and classification criteria for each subset.
4 Experiments
4.1 Setup
Models.
We tested SURGE on open-source and closed-source models of different sizes, including both chat models and code models. The closed-source models include GPT-4o (2024-08-06) (OpenAI, 2024b), GPT-4o-mini (2024-07-18) (OpenAI, 2024a), Claude-3.5-Sonnet (2024-10-22) (Anthropic, 2024a), and Qwen-Max (2025-01-25) (Team, 2024b). The open-source models include LLaMA-3.1-{8, 70}B-Instruct, LLaMA-3.3-70B-Instruct, Qwen-2.5-{0.5, 1.5, 3, 7, 14, 32, 72}B-Instruct, Qwen-2.5-Coder-{0.5, 1.5, 3, 7, 14, 32}B-Instruct and DeepSeek-V3 (671B).
Settings.
We tested the above models on SURGE under 3 settings: 0-shot w/o CoT, 0-shot w/ CoT, and few-shot w/ CoT. CoT here means whether we use Chain-of-Thought (Wei et al., 2022) prompting, allowing the models to think step by step, or ask the models to answer directly. We set the temperature to , i.e., employing greedy decoding.
| Model | Compiler | Standard | Optimization | ||||||
| GPT-4o | 14.75 | 14.21 | 22.08 | 15.33 | 14.61 | 20.02 | 7.66 | 1.63 | 1.22 |
4.2 Results
Table 2 presents the performance of 10 selected models across 8 sub-datasets of SURGE under 3 different settings. In this table, we provide a detailed breakdown of the models’ performance on the ML and BG sub-datasets. We present the complete results of all 21 models in Table 5 in Appendix A. From the results, several notable findings emerge:
SURGE demonstrates strong discriminative ability. Even the strongest models perform only moderately well, highlighting the value of our benchmark. The models exhibit significant performance differences across different subsets, reflecting the comprehensiveness of our dataset. Additionally, models that perform well in other tasks, such as Claude-3.5-Sonnet, also achieve substantial overall results on SURGE. This demonstrates the reasonableness of our dataset and its effectiveness in benchmarking LLMs as general-purpose surrogate code executors.
Chat models and coder models exhibit different performance patterns and are affected differently by prompting strategies. We observed that for chat and code models of the same size, code models outperform chat models in the zero-shot setting SURGE. However, in the other two settings, chat models perform better. This suggests that code models have stronger zero-shot surrogate capabilities, whereas chat models excel in reasoning and imitation abilities.
4.3 Anomaly Analysis
We conducted case studies on anomalous results to better understand model behavior.
Larger Models Underperforming. In some instances, larger models performed worse than their smaller counterparts. For example, on ML tasks, Qwen-2.5-14B-Instruct was outperformed by Qwen-2.5-0.5B-Instruct. Our case study revealed that while the 0.5B model’s reasoning was less detailed, it correctly grasped the code’s core logic. In contrast, the 14B model attempted more elaborate reasoning but often made critical misjudgments (e.g., in conditionals like ), leading to incorrect results. This suggests that on certain complex reasoning tasks, larger models may be prone to "overthinking" or following incorrect logical paths.
Zero-Shot Failures. We investigated cases of complete failure in the zero-shot setting, such as LLaMA-3.1-8B-Instruct scoring 0 on ML tasks for C++ and Rust, and on all FL tasks. For ML, the failures were attributed to the inherent difficulty of the test cases and specific language features (e.g., Rust’s strict type system). For FL tasks, the model incorrectly identified errors in all problems. For half of the correct problems, it wrongly predicted errors. For the other half that contained errors, it failed to predict the specific error messages accurately, leading to incorrect predictions in all cases.
5 Analysis
5.1 The Impact of Language Type
In the ML subset, LLMs perform best in Python, followed by C++. Python’s simple syntax makes it easier to process, while C++ benefits from its strong presence in programming and system-level tasks. Models also perform well in Julia, likely due to its clean syntax and similarity to Python. However, performance drops significantly in Rust, where strict ownership and lifetime rules introduce complexity, making it hard to predict.
In the BG subset, when predicting the output of buggy code, models excel in Python but struggle with C++. Python’s clear error messages aid prediction, whereas C++’s static typing, manual memory management, and potential for undefined behavior make error handling more difficult.
| Difficulty | 1 | 2 | 3 | 4 | 5 |
| Zero-shot | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Zero-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
| Few-shot Chain-of-Thought | |||||
| Claude-3.5-Sonnet | |||||
| GPT-4o | |||||
| LLaMA-3.1-8B-Instruct | |||||
| LLaMA-3.1-70B-Instruct | |||||
5.2 The Impact of Problem Difficulty
5.3 Prediction Accuracy & CPU Time
We explore the relationship between the execution time of the given code through running on CPUs and the accuracy of using LLMs as surrogate models to acquire the output. We categorize problems in TC into distinct bins based on their execution times and calculate the average accuracy for each model across all samples within the same bin, as shown in Figure 4. We observed the trend of prediction accuracy of the model falling as the actual execution time required for the corresponding program prolonged. It’s especially worth noting that for computational tasks that require execution time longer than 1 second, even state-of-the-art models struggle to obtain even one correct answer.
5.4 The Impact of Variable Type
The DR subset in SURGE examines model performance in predicting code behavior under different environmental factors. Specifically, DR includes variations in C++ compiler version, C++ standard, and compilation optimization settings. Table 3 presents model performance across different prompting strategies in DR.
5.5 In-depth Look at Model Behavior
We looked into LLMs’ surrogation process and summarized three key observations:
Deep Semantic and Logical Reasoning. LLMs demonstrate a strong ability to interpret complex, language-specific semantics. For instance, when analyzing C++ code with bit fields, models correctly reasoned about the effects of signed vs. unsigned types, bit-width limitations, and how value overflows are handled under two’s complement representation. This indicates an understanding of low-level computational logic.
Stateful Algorithmic Tracing. For algorithmic tasks, LLMs effectively simulate the execution process by tracing the program’s state. The models’ chain-of-thought process shows them iterating through loops, updating variables step-by-step, and applying conditional logic as an interpreter would.
Identification of Computational Boundaries. In a brute-force Traveling Salesman Problem (TSP) solver, models could correctly understand the overall goal and algorithm, but often failed to predict the precise final floating-point value. This suggests their "execution" is a form of sophisticated logical inference, which falters when faced with complex combinatorial calculations that exceed their internal reasoning capacity.
5.6 Error Analysis
To further understand the model’s performance and limitations regarding serving as general surrogate models, we categorize the Errors made by Claude-3.5-Sonnet, GPT-4o, and Llama-3.1-8B-Instruct and Llama-3.1-70B-Instructin the CL subset of SURGE.
We first combined LLM-assisted annotation and manual verification to identify most typical error types: 1. Code Language Knowledge: evaluating foundational programming language proficiency, 2. Context Awareness: measuring understanding of long text and repository-level code, 3. Conceptual Understanding: assessing comprehension of programming concepts, 4. Calculation Process: verifying computational step accuracy, 5. Code Understanding: testing comprehension of code logic and structure, 6. Data Interpretation: evaluating data processing and analysis capabilities, and 7. Calculation Accuracy: Measuring precision in scientific computations. We then use LLM to label errors in the model’s responses. The categorized error statistics are demonstrated in Figure 5.
As models prompted with CoT exhibit fewer instances of most error types compared to zero-shot, especially for the Code Understanding capability of Llama. The main error types are Code Understanding, Calculation Process, and Calculation Accuracy. For zero-shot, the primary error is accuracy, but for CoT, the most frequent error is Calculation Process. This suggests that CoT can better grasp the overall code logic and produce more correct results, but it may still make mistakes in the chain of thought process. In general, CoT has fewer and smaller errors. From the model perspective, Llama has a clear lead in Conceptual Understanding errors, indicating its weaker ability to understand concepts.
5.7 Training Scale Analysis
We also investigate whether training scaling can affect LLMs’ surrogate execution capabilities on the FL task. We trained models of varying sizes on different amounts of training data sampled from the same distribution, and tested them to predict the error feedbacks for incorrect proofs.
As demonstrated in Figure 2, both model size and training steps are crucial factors in determining surrogate execution accuracy. As we scale from 0.5B to 7B parameters, models consistently show improved learning efficiency and higher performance ceilings throughout the training process. Larger models learn faster in the early stages and continue to improve for longer before plateauing, suggesting better utilization of the training data.
These empirical observations align with established scaling laws in language modeling, indicating that surrogate execution capabilities follow similar scaling patterns as other language tasks.
6 Conclusion
We introduce SURGE, a holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. Through extensive empirical study, we argue that there remains significant room for further improvements on grounding LLMs to facilitate general surrogate models. Our findings not only chart the current landscape but also illuminate a clear path for future research.
Acknowledgements
This work was independently conducted by the authors and is self-funded by BL without institutional affiliation.
Limitations
Despite its comprehensive evaluation, our study has several limitations. LLMs remain approximators rather than exact code executors, often struggling with edge cases, intricate runtime behaviors, and execution-dependent state changes. While SURGE covers diverse execution scenarios, it does not encompass all specialized environments, such as hardware-dependent simulations or real-time systems. Additionally, LLMs may generate plausible but incorrect outputs, particularly in complex logical dependencies or undefined behaviors, making error detection challenging. Our scaling study is constrained by computational resources, limiting the assessment of extremely large models or extensive training data distributions. Furthermore, security risks remain, as LLMs may fail to recognize vulnerabilities, potentially misjudging harmful code. Finally, our benchmark operates in a controlled setting, whereas real-world software development involves dynamic interactions and iterative debugging, which are not fully captured in our study. Future work should focus on improving LLMs’ reasoning abilities, enhancing robustness in execution prediction, and integrating them with traditional program analysis techniques for practical deployment.
Appendix
Appendix A Complete Main Results
@@
Appendix B Prompts
B.1 Prompts for Dataset Refactoring
ML:
CL:
BG:
Appendix C Details of SURGE
C.1 ML
| Java | C# | Rust | Julia | Python | C++ | C |
| 25 | 20 | 20 | 26 | 18 | 21 | 20 |
In ML, the usage distribution of various programming languages is shown in Table 6. We selected a variety of languages, including Java, C#, Rust, Julia, Python, C++, and C, to evaluate the model’s ability to handle multilingual code. This diverse selection helps to comprehensively assess the model’s performance across different languages.
C.1.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.1.2 Demo Questions
C.2 CL
| Python | C++ | JavaScript |
| 50 | 51 | 49 |
| Difficulty | JavaScript | CPP | Python |
| 1 | 10 | 11 | 11 |
| 2 | 6 | 4 | 6 |
| 3 | 12 | 14 | 12 |
| 4 | 8 | 8 | 9 |
| 5 | 13 | 14 | 12 |
In CL, we selected competition problems of varying difficulty, each with solutions in Python, C++, and JavaScript. You can see the distribution of language in Table 7, and the distribution of problem difficulty in Table 8. This selection allows us to test the model’s cross-language capabilities and its ability to handle problems of different difficulty levels.
C.2.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.2.2 Demo Questions
C.3 RL
| Python | C++ |
| 24 | 36 |
In RL, the distribution of programming language usage is shown in Table 9. We utilized five GitHub repositories for this study, consisting of two Python projects and three C++ projects. Each repository contains a set of ten or more test cases, providing a diverse set of data for evaluation across different programming languages.
C.3.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.3.2 Demo Questions
C.4 SC
C.4.1 Tasks Descriptions
The scientific computing component of SURGE consists of 4 carefully curated areas, aiming to evaluate model performance on computational tasks that exhibit a time-consuming nature as well as applicational values in scientific computing areas. In this section, we provide a detailed description of each component.
Numerical Optimization.
In this task, the model is given a program that solves an optimization problem through gradient descent. The query may be the optimized value (min) or the optimal point (argmin). We carefully select four functions, which consist of: a simple quadratic function, Rosenbrock Function, Himmelblau’s Function, and a polynomial function with linear constraints. For each function, we will select multiple different hyperparameter configurations to assess the model’s performance. These four functions provide a systematic evaluation of the model’s potential to serve as a surrogate model in this field. As the quadratic function is solvable without need the to run the gradient descent, the model may solve it through world knowledge. The Rosenbrock function is known for its narrow, curved valley containing the global minimum, making it difficult for optimization algorithms to converge. Therefore the output is highly dependent on hyperparameters (initial point, learning rate, maximum steps), thus the model must execute code in its reasoning process to acquire the answer. Himmelblau’s function has multiple local minima, also posing sensitivity to hyperparameters.
PDE Solving.
We consider three types of Partial Differential Equations: the 1D Heat Equation, the 2D Wave Equation, and the 2D Laplace Equation. For the 1D Heat Equation, we focus on solving the following equation:
For the 2D Laplace Equation, we aim to solve the equation:
Lastly, for the 2D Wave Equation, we work on solving the following equation:
We solve 1D Heat Equation and 2D Wave Equation using the Explicit Finite Difference Method. For the 2D Laplace Equation, we solve it using the Gauss-Seidel Method. The model is then queried on the values of and .
Fourier Transform (FFT)
We implement FFT using the Cooley-Tukey Algorithm and query the model to give the magnitude of the top 10 values.
ODE Solving
For solving ordinary differential equations, we constructed three different equations and implemented the Euler Method and the Runge-Kutta Method so solve these equations.
C.4.2 Evaluation Metrics
Relative Absolute Error (RAE).
Given a scalar ground truth value and a model prediction , the Relative Absolute Error (RAE) is defined as:
For cases involving multiple entries, such as tensors or vectors, the following alignment procedure is applied: (1) if the prediction contains fewer elements than the ground truth, the prediction is padded with zeros until it matches the length of the ground truth; (2) if the prediction has more elements than the ground truth, it is truncated to match the ground truth length. The average RAE is then computed by averaging the RAE for each corresponding element.
Exact Matching.
For tasks involving position-based predictions, such as binary search, we adapt exact matching, as the accuracy of the algorithm is determined by comparing the exactness of the estimated result to the true result. This evaluation method checks if the estimated solution matches the ground truth exactly, typically using string or sequence matching. For such tasks, an exact match is considered a success, and any discrepancy between the ground truth and the estimate results in failure. Formally, given a string and the model’s prediction , the Exact Matching is given by:
where is the indicator function.
C.4.3 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.4.4 Demo Questions
C.5 TC
C.5.1 Tasks Descriptions of Time Consuming (TC)
The time consuming component of SURGE is comprised of 4 tasks in for computationally expensive areas, covering a spectrum of Linear Algebra, Sorting, Searching, Monte Carlo Simulations and String Matching Programs. Some of these tasks take hours to complete, showing their potential to benchmark LLM’s ability to reason through lengthy computations.
Linear Algebra.
In this task, we are focused on acquiring key properties in linear algebra given square matrices of varying sizes. In particular, we query the model on solving LU decomposition, QR decomposition, the largest eigenvalue and eigenvector using the power method, and the inversion matrix.
Sorting And Searching.
We include four classical algorithmic problems in this area, namely Hamiltonian Cycle, Traveling Salesman Problem (TSP), Sorting an array of real numbers and Searching. For Hamiltonian Cycle, we adopt the backtracking algorithm. Specifically, we randomly generate graphs with vertices from 4 to 100 and ask the model to find whether a Hamiltonian cycle exists. For TSP, we implement a naive brute-force algorithm and ask the model to find the length of the optimal path. For Sorting, we adopt the bubble sort, quick sort, and merge sort algorithms. For each algorithm, we consider different list sizes from 5 to 100 and generate 10 test cases for each list size. The evaluation metric is the rank correlation (also Spearman’s ). Lastly, for searching, we adopt binary search and query the model on randomly generated lists of varying sizes.
Monte Carlo Estimation.
We adopt Monte Carlo simulation to estimate the values of specific real numbers (e.g. ), as well as a future stock price prediction that follows the Brownian motion. We alter the number of samples used in Monte Carlo estimation, resulting in varying program outcomes.
String Matching Program.
We adopt the naive string matching, KMP, and Rabin-Karp algorithms. For each algorithm, we randomly generate text and pattern with varying lengths, and query the model on the existence and position of the matching.
C.5.2 Evaluation Metrics
Rank Correlation.
Rank Correlation (Spearman, 1904), also referred to as Spearman’s , is used to assess sorting tasks by measuring the correlation between the estimated ordinal ranking and the ground truth, which can be written as:
where and denote the true and estimated rankings, respectively, and Cov and represent the covariance and standard deviation of the respective sequences.
C.5.3 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.5.4 Demo Questions
C.6 BG
| Java | Python | C++ |
| 51 | 45 | 54 |
| Error Type | Java | Python3 | CPP |
| == and = confusion | 5 | 6 | 5 |
| undefined keywords | 6 | 3 | 5 |
| parentheses mismatch | 5 | 5 | 6 |
| indexing error | 10 | 9 | 11 |
| undefined objects | 11 | 9 | 8 |
| unclosed string | 7 | 5 | 7 |
| conditional statement error | 10 | 8 | 9 |
| undefined methods | 8 | 3 | 6 |
| colon missing | 5 | 7 | 8 |
| wrong comment mark | 9 | 1 | 9 |
| variable value error | 2 | 2 | 4 |
| operation error | 2 | 2 | 3 |
| other error | 4 | 2 | 1 |
| statement separation | 4 | 0 | 7 |
| indentation error | 0 | 4 | 0 |
| Double Bugs | 10 | 8 | 10 |
| Triple Bugs | 12 | 10 | 11 |
| Quadruple Bugs | 8 | 5 | 9 |
In BG, the distribution of language usage across categories is shown in Table 10, indicating a balanced usage of Java, Python, and C++. Table 11 presents a detailed breakdown of bug types and their frequency across different languages. This distribution allows us to assess the model’s ability to handle a variety of bugs across multiple programming languages.
C.6.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.6.2 Demo Questions
C.7 DR
C.7.1 System Prompts
Zero-shot Chain-of-Thought:
Zero-shot:
Few-shot Chain-of-Thought:
C.7.2 Demo Questions
C.8 FL
Hard for Latex to Compile, check the source code.
Appendix D Training Details
For training, we employ Llama-Factory (Zheng et al., 2024) as the LLM training platform. Table 12 shows our training hyperparameters.
| Parameter | Value |
| Train batch size | 128 |
| Learning rate | 1.0e-5 |
| Number of epochs | 2.0 |
| LR scheduler | cosine |
| Warmup ratio | 0.1 |
| Precision | bf16 |